OpenStudy (anonymous):

If sin theta = (sqrt 3)/2, which could not be the value of theta? A. 60 degrees B. 120 degrees C. 240 degrees D. 420 degrees I know that the solutions for theta are pi/3 and 2pi/3, but how does that correlate?

OpenStudy (anonymous):

\[\sin ^{-1}\left(\frac{\sqrt{3}}{2}\right)=\frac{\pi }{3}=60 {}^{\circ} \]

OpenStudy (anonymous):

Thanks! I was almost 100% sure this had something to do with the unit circle, but I wasn't quite sure.

OpenStudy (anonymous):

Thank you for the medal.